# Contrastive Learning and Adversarial Disentanglement for Privacy-Preserving Task-Oriented Semantic Communications

### v1

@@ Abstract
Task-oriented semantic communication systems have emerged as a promising approach to achieving efficient and intelligent data transmission, where only information relevant to a specific task is communicated. However, existing methods struggle to fully disentangle task-relevant and task-irrelevant information, leading to privacy concerns and subpar performance. To address this, we propose an information-bottleneck method, named CLAD (contrastive learning and adversarial disentanglement). CLAD leverages contrastive learning to effectively capture task-relevant features while employing adversarial disentanglement to discard task-irrelevant information. Additionally, due to the lack of reliable and reproducible methods to gain insight into the informativeness and minimality of the encoded feature vectors, we introduce a new technique to compute the information retention index (IRI), a comparative metric used as a proxy for the mutual information between the encoded features and the input, reflecting the minimality of the encoded features. The IRI quantifies the minimality and informativeness of the encoded feature vectors across different task-oriented communication techniques. Our extensive experiments demonstrate that CLAD outperforms state-of-the-art baselines in terms of task performance, privacy preservation, and IRI. CLAD achieves a predictive performance improvement of around 2.5-3%, along with a 77-90% reduction in IRI and a 57-76% decrease in adversarial accuracy.
@@
Contrastive learning, disentangled representation learning, information-bottleneck, task-oriented communications, semantic communications.
@@ I Introduction
With the growing success and popularity of deep learning (DL) in various wireless communication applications [11], many emerging task-oriented communication systems have adopted DL approaches to encode task-relevant information to improve task performance and efficiency of the communication system[12, 13, 14]. Nevertheless, most proposed schemes do not focus quantifying or benchmarking the amount of information that the encoded features retain about the input, primarily due to the computational difficulty of estimating mutual information and the lack of a unified methodology that provides fair and reproducible results. This omission is critical, as understanding the mutual information is key to developing systems with improved bandwidth and latency. Moreover, there has been little emphasis on explicitly discarding task-irrelevant information, often leading to a trade-off between task performance and information compression.
@@
  • We evaluate our proposed task-oriented communication system in diverse channel conditions. It is tested against several existing task-oriented communication methods, demonstrating improved task performance, enhanced privacy preservation, and reduced amount of stored information across a wide range of transmission scenarios.
@@ IV Problem Description
In this section, we identify the primary challenges that arise when transmitting features over a communication channel and utilizing them for a downstream classification task. Our goal is to ensure that the transmitted features contain only the minimum necessary information required for the downstream task to preserve privacy and maximize efficiency. Furthermore, we argue that it is necessary to have a fair, reproducible, and unified method to obtain comparative values that act as a proxy for mutual information between the encoded features and the input data to allow effective benchmarking of different task-oriented communication systems.
@@ IV-A Minimum Necessary Information
Following [30], the Minimum Necessary Information (MNI) criterion for an ideal representation $\hat{Z}$ must satisfy the following key principles:
@@
Given the above we conclude that in an optimal case we must have
@@ IV-D Limitations of Variational Information Bottleneck (VIB)
where $I(\hat{Z};X)$ measures the amount of information retained from the input $X$, and $I(\hat{Z};Y)$ represents the informativeness of $Z$ for predicting $Y$. Hyperparameter $\beta$ controls the trade-off between preserving task-relevant information and discarding irrelevant information. To maximize $I(\hat{Z};Y)$, a cross-entropy based loss is used, and the Kullback–Leibler divergence is used to to minimize $I(\hat{Z};X)$ [15, 12].
@@
  • Difficulty in tuning $\beta$: The choice of $\beta$ significantly impacts performance, with an improper setting leading to either poor classification accuracy, or leakage of unnecessary details. Finding the optimal $\beta$ for each task is in-practice non-trivial.
@@
To address the limitations of VIB, we propose two key modifications. Firstly, we use supervised contrastive learning as shown in Fig. 2 to directly optimize $I(\hat{Z};Y)$, thereby improving predictive performance. Secondly, we introduce adversarial methods to disentangle task-relevant and task-irrelevant information, ensuring that we approach the MNI while maintaining classification accuracy and discarding unnecessary or private details and minimizing the reliance on the trade-off between $I(\hat{Z};X)$ and $I(\hat{Z};Y)$ through a hyperparameter $\beta$.
@@ V Proposed Method: CLAD
We propose CLAD, a comprehensive task-oriented communication scheme designed to achieve strong task performance and improved privacy, through effective disentanglement of task-relevant and task-irrelevant information. The method utilizes two key encoders: the TRE, which maps the input into the task-relevant channel codeword $Z_{1}$, and the task-irrelevant encoder (TIE), which maps the input into task-irrelevant channel codeword $Z_{2}$. The task performance is optimized through contrastive learning, which aims to maximize the mutual information $I(\hat{Z_{1}};Y)$, ensuring that $\hat{Z_{1}}$ captures the most informative features for downstream classification. The disentanglement is achieved through reconstruction learning to capture task-irrelevant information in $\hat{Z_{2}}$, and adversarial training is utilized to minimize the mutual information $I(\hat{Z_{1}};\hat{Z_{2})}$, thus promoting independence between the two feature representations. The different components and training stages of CLAD are visualized in Fig. 3. These components are optimized together to ensure both high task accuracy and effective disentanglement of information, corresponding to the following maximization objective:
@@
(a) Stage 1: Learning task-irrelevant features.
(b) Stage 2: Learning task-relevant information and disentangling task-irrelevant information.
(c) Stage 3: Classifier training.
Figure 3: Three stages for training CLAD.
@@ V-A Contrastive Loss for Task-Relevant Features
$\mathcal{L}_{\text{contrast}}=\frac{1}{\sum_{i\neq j}\mathds{1}_{y_{i}=y_{j}}}\sum_{i\neq j}\mathds{1}_{y_{i}=y_{j}}\log S_{ij},$ (12)
@@
and thus minimizing $\mathcal{L}_{\text{contrast}}$ maximizes a lower bound of $I(\hat{Z},Y)$. Increasing $B$ raises $\log(B-1)$, thereby strengthening this lower bound and enhancing performance by preserving more task-relevant information.
@@ V-F Information Retention Index Across Different Methods
  • All reconstructors are trained with the settings and hyperparameters for the same number of epochs.
@@ VI Experimental Evaluations and Discussion
In this section, we present the experimental setup used to evaluate CLAD. We start by describing the datasets used in our experiments, followed by a discussion of the baseline methods, neural architectures, and the experimental setup. Finally, we present detailed evaluations and analysis of the results. For brevity, we omit results on other datasets, such as CIFAR-10; however, these can be found in the accompanying code repository [^1] The source code, models and results are available at https://github.com/OmarErak/CLAD.
@@ VI-A1 Datasets
The Colored MNIST and Colored FashionMNIST datasets are extensions of the standard MNIST [35] and FashionMNIST [36] datasets, each consisting of 60,000 28x28 grayscale images. In Colored MNIST, handwritten digits (0-9) are overlaid on colored backgrounds, while in Colored FashionMNIST, clothing items from 10 categories (e.g., T-shirts, coats, shoes) are similarly displayed on colored backgrounds. The introduction of background colors adds additional task-irrelevant information, creating a more challenging setup for the model to disentangle task-relevant features relevant to classifying digits or clothing items from background-related attributes. Furthermore, incorporating background color labels enables the evaluation of attribute inference attacks, where an adversary is trained to predict background color. This setup provides insight into the model’s ability to protect against such attacks while maintaining disentanglement between task-relevant and task-irrelevant features.
@@ VI-A2 Neural Network Architectures
In this work, we employ a deep neural network (DNN) architecture for the classification tasks across both Colored MNIST and Colored FashionMNIST datasets. The architecture consists of convolutional and fully connected layers, with a latent dimension $d=64$. This architecture, detailed in Table I, serves as the backbone for the task-relevant encoder (transmitter) and the downstream classification task (receiver). The same architecture is used across all baselines to ensure a consistent and fair comparison of methods. To compute the IRI, for each method we train multiple reconstructors that aim to reconstruct the original input $\boldsymbol{x}$ from the latent representation $\hat{\boldsymbol{z}}$. The reconstructor architecture is described in Table II, and it is used consistently across all methods for IRI computations.
@@ VI-A3 Channel Conditions
TABLE I: DNN Structure for the transmitter (encoder) and the receiver (classifier)
@@
TABLE II: The architecture settings for the reconstructors used to evaluate IRI
@@ VI-B2 Variational Information Bottleneck (VIB)
The variational information bottleneck (VIB) framework aims to compress the input $X$ into latent representation $Z$ while retaining sufficient information for predicting $Y$. VIB balances $I(Z;X)$ and $I(Z;Y)$ via hyperparameter $\beta$. In our experiments, we provide results for different values of $\beta$ to illustrate how varying the trade-off between compression and task relevance impacts task performance, IRI, and privacy.
@@ VI-C Evaluation Metrics
To evaluate the effectiveness of CLAD, we utilize three key metrics: task performance through classification accuracy, IRI, and attribute inference attack accuracy. These metrics provide insights into how well the method performs in terms of classification accuracy, information retention, and privacy preservation. Additionally, we assess all methods at different channel SNRs to measure performance under various dynamic transmission conditions.
@@ VI-C2 Information Retenetion Index (IRI)
Figure 4: Accuracy at different SNRs for (a) the Colored MNIST dataset, and (b) the Colored FashionMNIST dataset.
@@ VI-C3 Attribute Inference Attack
TABLE III: Evaluation of different methods on Colored MNIST dataset at SNR = 12 dB.
@@
(a) Colored MNIST Reconstructions
(b) Colored Fashion MNIST Reconstructions
Figure 6: Original and reconstructed images using the encoded feature vectors from different methods.
@@ VI-D Results and Analysis
In this subsection we thoroughly analyze and discuss the performance of the proposed method CLAD against the two baselines, DeepJSCC and VIB, on both Colored MNIST and Colored FashionMNIST datasets.
TABLE IV: Evaluation of different methods on Colored FashionMNIST dataset at SNR = 12 dB.
@@ VI-D1 Colored MNIST Results
Fig.  4(a) shows the accuracy as a function of the SNR for each method. Here, it is evident that CLAD outperforms all the baselines at different channel conditions. This is a strong indication that maximizing $I(\hat{Z};Y)$ through contrastive learning yields better feature extraction and classification performance, even with undesirable and high noise conditions. As seen in Table III, CLAD outperforms both DeepJSCC and VIB, achieving the highest classification accuracy of 98.42% at SNR = 12 dB. This represents a 0.46% improvement over DeepJSCC (97.96%) and a 0.52% improvement over VIB at $\beta=0.001$ (97.90%). The higher accuracy of CLAD can be attributed to its ability to learn a more disentangled and task-relevant latent space using contrastive learning and adversarial disentanglement.
From Table III, CLAD achieves a very low IRI value of 0.039, drastically outperforming DeepJSCC (0.608) and VIB at $\beta=0.001$ (0.1931). This substantial difference shows that CLAD is effective at disentangling and removing task-irrelevant information from its encoded representations. This is further supported by the significantly lower adversarial accuracy of CLAD compared to DeepJSCC and VIB at $\beta=0.001$. From Fig. 6(a), we can visually see that image reconstructions from CLAD excel in retaining task-relevant information and preserving the core structure of the digits while discarding irrelevant information such as background color and other style features such as thickness or slant in the digit. This contrasts with DeepJSCCC and VIB, which retain unnecessary information such as background color and thickness and slant of the digit.
Interestingly, from Table III, VIB at lower $\beta$ values (e.g., $\beta=0.01$ and $\beta=0.1$) achieves IRI values that are closer to CLAD. However, the major drawback of VIB is that, while it improves privacy, informativeness and minimality, it comes at the cost of accuracy, especially as $\beta$ increases, as shown in the table. For example, at $\beta=0.1$, VIB achieves even lower IRI, however its classification accuracy drops to 93.29%, which is significantly lower than both CLAD and DeepJSCC. This demonstrates that CLAD maintains task performance while better preserving accuracy compared to VIB.

### v2

# Contrastive Learning and Adversarial Disentanglement for Task-Oriented Semantic Communications

@@ Abstract
Task-oriented semantic communication systems have emerged as a promising approach to achieving efficient and intelligent data transmission, where only information relevant to a specific task is communicated. However, existing methods struggle to fully disentangle task-relevant and task-irrelevant information, leading to privacy concerns and subpar performance. To address this, we propose an information-bottleneck method, named CLAD (contrastive learning and adversarial disentanglement). CLAD utilizes contrastive learning to effectively capture task-relevant features while employing adversarial disentanglement to discard task-irrelevant information. Additionally, due to the lack of reliable and reproducible methods to gain insight into the informativeness and minimality of the encoded feature vectors, we introduce a new technique to compute the information retention index (IRI), a comparative metric used as a proxy for the mutual information between the encoded features and the input, reflecting the minimality of the encoded features. The IRI quantifies the minimality and informativeness of the encoded feature vectors across different task-oriented communication techniques. Our extensive experiments demonstrate that CLAD outperforms state-of-the-art baselines in terms of semantic extraction, task performance, privacy preservation, and IRI. CLAD achieves a predictive performance improvement of around 2.5-3%, along with a 77-90% reduction in IRI and a 57-76% decrease in adversarial attribute inference attack accuracy.
@@
Contrastive learning, disentangled representation learning, information-bottleneck, semantic communications, task-oriented communications.
@@ I Introduction
With the growing success and popularity of deep learning (DL) in various wireless communication applications [11, 12], many emerging task-oriented communication systems have adopted DL approaches to encode task-relevant information to improve task performance and efficiency of the communication system[13, 14, 15]. Nevertheless, most proposed schemes do not focus quantifying or benchmarking the amount of information that the encoded features retain about the input, primarily due to the computational difficulty of estimating mutual information and the lack of a unified methodology that provides fair and reproducible results. This omission is critical, as understanding the mutual information is key to developing systems with improved bandwidth and latency. Moreover, there has been little emphasis on explicitly discarding task-irrelevant information, often leading to a trade-off between task performance and information compression.
@@
  • We evaluate our proposed task-oriented communication system in diverse channel conditions. It is tested against several existing task-oriented communication methods, demonstrating improved task performance, enhanced privacy awareness, and reduced amount of stored information across a wide range of transmission scenarios.
@@ IV Problem Description
In this section, we identify the primary challenges that arise when transmitting features over a communication channel and utilizing them for a downstream classification task. Our goal is to ensure that the transmitted features contain only the minimum necessary information required for the downstream task to maximize efficiency whilst being privacy aware. Furthermore, we argue that it is necessary to have a fair, reproducible, and unified method to obtain comparative values that act as a proxy for mutual information between the encoded features and the input data to allow effective benchmarking of different task-oriented communication systems.
@@ IV-A Minimum Necessary Information
Following [31], the Minimum Necessary Information (MNI) criterion for an ideal representation $\hat{Z}$ under ideal transmission conditions must satisfy the following key principles:
@@
Given the above we conclude that in an optimal case under ideal channel conditions we must have
@@ IV-D Limitations of Variational Information Bottleneck (VIB)
where $I(\hat{Z};X)$ measures the amount of information retained from the input $X$, and $I(\hat{Z};Y)$ represents the informativeness of $Z$ for predicting $Y$. Hyperparameter $\beta$ controls the trade-off between preserving task-relevant information and discarding irrelevant information. To maximize $I(\hat{Z};Y)$, a cross-entropy based loss is used, and the Kullback–Leibler divergence is used to to minimize $I(\hat{Z};X)$ [16, 13, 15].
@@ V Proposed Method: CLAD
The core idea of CLAD is to create an end-to-end communication system that explicitly learns to disentangle task-relevant information from task-irrelevant content, enabling both high task accuracy and improved privacy. Rather than simply combining existing techniques, CLAD integrates supervised contrastive learning, adversarial training, and reconstruction-based supervision into a unified, structured framework that aligns with the Minimum Necessary Information (MNI) principle.
Each component in CLAD is designed to target a specific challenge:
  • Contrastive learning maximizes the informativeness of the representation with respect to the downstream task.
  • Adversarial disentanglement promotes independence between task-relevant and irrelevant features, suppressing information leakage.
  • Reconstruction learning ensures that task-irrelevant features are adequately captured.
To ensure effective optimization, CLAD adopts a three-stage training strategy, where each stage isolates and refines a different component of the system. This design is intentional — it prevents interference between competing objectives, stabilizes training, and encourages modular reuse of the encoders across stages.
CLAD utilizes two key encoders: the TRE, which maps the input into the task-relevant channel codeword $Z_{1}$, and the task-irrelevant encoder (TIE), which maps the input into task-irrelevant channel codeword $Z_{2}$. The task performance is optimized through contrastive learning, which aims to maximize the mutual information $I(\hat{Z_{1}};Y)$, ensuring that $\hat{Z_{1}}$ captures the most informative features for downstream classification. The disentanglement is achieved through reconstruction learning to capture task-irrelevant information in $\hat{Z_{2}}$, and adversarial training is utilized to minimize the mutual information $I(\hat{Z_{1}};\hat{Z_{2})}$, thus promoting independence between the two feature representations. The different components and training stages of CLAD are visualized in Fig. 3. These components are optimized together to ensure both high task accuracy and effective disentanglement of information, corresponding to the following maximization objective:
@@
Figure 3: Three stages for training CLAD
@@ V-A Contrastive Loss for Task-Relevant Features
$\mathcal{L}_{\text{contrast}}=-\frac{1}{\sum_{i\neq j}\mathds{1}_{y_{i}=y_{j}}}\sum_{i\neq j}\mathds{1}_{y_{i}=y_{j}}\log S_{ij},$ (12)
@@
and thus minimizing $\mathcal{L}_{\text{contrast}}$ maximizes a lower bound of $I(\hat{Z},Y)$. Increasing $B$ raises $\log(B-1)$, thereby strengthening this lower bound and enhancing performance by preserving more task-relevant information. Although the derived bound can be loose with a small number of negative samples [17], we mitigate this by sampling large batches (2048) during contrastive training.
@@ V-E Training Procedure
The separation of training stages is for stability, and also reflects a modular decomposition of the CLAD objective. Each stage isolates the gradient flow to the component most relevant for that term, making the multi-stage training strategy a practical approximation to optimizing $\mathcal{L}_{\text{CLAD}}$ holistically.
@@ V-F Information Retention Index Across Different Methods
  • All reconstructors are trained with the same settings and hyperparameters for the same number of epochs.
@@
While methods such as MINE [36] offer tighter approximations of MI, they require adversarial or variational training, are sensitive to hyperparameters, and are not directly comparable across model architectures. In contrast, IRI allows for consistent, fair benchmarking with minimal assumptions and training overhead. We view IRI as a relative measure, similar to a diagnostic tool, that enables empirical analysis of minimality and informativeness in task-oriented systems. We note that while SSIM is effective for evaluating visual similarity in image-based applications, it is inherently limited to image data. For non-visual modalities such as text or tabular data, alternative reconstruction-based metrics (e.g., BLEU score) can be adopted to extend the IRI framework to broader task domains.
@@ VI Experimental Evaluations and Discussion
In this section, we present the experimental setup used to evaluate CLAD. We start by describing the datasets used in our experiments, followed by a discussion of the baseline methods, neural architectures, and the experimental setup. Finally, we present detailed evaluations and analysis of the results. [^1] The source code, models and results are available at https://github.com/OmarErak/CLAD.
@@ VI-A1 Datasets
The Colored MNIST and Colored FashionMNIST datasets are extensions of the standard MNIST [37] and FashionMNIST [38] datasets, each consisting of 60,000 28x28 grayscale images. In Colored MNIST, handwritten digits (0-9) are overlaid on colored backgrounds, while in Colored FashionMNIST, clothing items from 10 categories (e.g., T-shirts, coats, shoes) are similarly displayed on colored backgrounds. The introduction of background colors adds additional task-irrelevant information, creating a more challenging setup for the model to disentangle task-relevant features relevant to classifying digits or clothing items from background-related attributes. Furthermore, incorporating background color labels enables the evaluation of attribute inference attacks, where an adversary is trained to predict background color. This setup provides insight into the model’s ability to protect against such attacks while maintaining disentanglement between task-relevant and task-irrelevant features. To further evaluate the scalability and robustness of the proposed method on more complex and natural datasets, we also incorporate CIFAR-10 [krizhevsky2009learningmultiple]. CIFAR-10 consists of 60,000 32x32 color images in 10 classes, including airplanes, cars, birds, cats, and other natural objects. Compared to MNIST-based datasets, CIFAR-10 introduces significantly more visual variability and semantic richness, making it a more challenging benchmark.
@@ VI-A2 Neural Network Architectures
In this work, we employ a deep neural network (DNN) architecture for the classification tasks across all datasets. These architectures consists of convolutional and fully connected layers, with a latent dimension $d=64$. The architectures, detailed in Table I and Table III, serve as the backbone for the task-relevant encoder (transmitter) and the downstream classification task (receiver). The same architecture is used across all baselines to ensure a consistent and fair comparison of methods. To compute the IRI, for each method we train multiple reconstructors that aim to reconstruct the original input $\boldsymbol{x}$ from the latent representation $\hat{\boldsymbol{z}}$. The reconstructor architectures are described in Table II and Table IV, and it is used consistently across all methods for IRI computations.
@@ VI-A3 Channel Conditions
TABLE I: DNN Structure for the transmitter (encoder) and the receiver (classifier) used for Colored MNIST and FashionMNIST
@@
TABLE II: Architecture settings for the reconstructors used with Colored MNIST and FashionMNIST to evaluate IRI
@@
TABLE III: DNN structure for the transmitter (encoder) and the receiver (classifier) used for CIFAR-10
Layer
Output
Dimensions
Transmitter
Conv + ReLU $\times$2
ResNet Block
Conv + ReLU $\times$2
Reshape + FC + Tanh
128$\times$32$\times$32
128$\times$16$\times$16
4$\times$4$\times$4
$d$
Receiver
FC + ReLU + Reshape
Conv + ReLU $\times$2
ResNet Block
Pooling Layer
FC + Softmax
64$\times$4$\times$4
512$\times$4$\times$4
512$\times$4$\times$4
512
10
TABLE IV: Architecture settings for the reconstructors used with CIFAR-10 to evaluate IRI
Layer name
Output
dimensions
Reconstructor Fully Connected (FC) + ReLU 512$\times$4$\times$4
Deconv (512$\rightarrow$256) + ReLU 256$\times$8$\times$8
Deconv (256$\rightarrow$128) + ReLU 128$\times$16$\times$16
Deconv (128$\rightarrow$64) + ReLU 64$\times$32$\times$32
Conv (64$\rightarrow$3) + Sigmoid 3$\times$32$\times$32
@@ VI-B2 Variational Information Bottleneck (VIB)
The variational information bottleneck (VIB) framework aims to compress the input $X$ into latent representation $Z$ while retaining sufficient information for predicting $Y$. VIB balances $I(Z;X)$ and $I(Z;Y)$ via hyperparameter $\beta$. In our experiments, we provide results for different values of $\beta$ to illustrate how varying the trade-off between compression and task relevance impacts task performance, IRI, and privacy. The work in [13] is based on VIB, and therefore, we reproduce their results as a baseline comparison.
@@ VI-C Evaluation Metrics
To evaluate the effectiveness of CLAD, we utilize three key metrics: task performance through classification accuracy, IRI, and attribute inference attack accuracy. These metrics provide insights into how well the method performs in terms of classification accuracy, information retention, and privacy awareness. Additionally, we assess all methods at different channel SNRs to measure performance under various dynamic transmission conditions.
@@ VI-C2 Information Retenetion Index (IRI)
Figure 4: Accuracy at different SNRs for (a) the Colored MNIST dataset, (b) the Colored FashionMNIST dataset and (c) the CIFAR-10 dataset .
@@ VI-C3 Attribute Inference Attack
TABLE V: Evaluation of different methods on Colored MNIST dataset at SNR = 12 dB.
@@ VI-D Results and Analysis
In this subsection we thoroughly analyze and discuss the performance of the proposed method CLAD against the two baselines, DeepJSCC and VIB, on all three aforementioned datasets.
TABLE VI: Evaluation of different methods on Colored FashionMNIST dataset at SNR = 12 dB.
@@
TABLE VII: Evaluation of different methods on CIFAR10 dataset at SNR = 12 dB.
Method
Classification
Accuracy (%)
IRI
DeepJSCC 91.21 0.2043
VIB (Beta=0.001) 91.49 0.1132
VIB (Beta=0.01) 86.42 0.0456
VIB (Beta=0.1) 78.31 0.0321
VIB (Beta=1) 9.67 0.0092
CLAD (Ours) 92.33 0.0471
@@ VI-D1 Colored MNIST Results
Fig.  4(a) shows the accuracy as a function of the SNR for each method. Here, it is evident that CLAD outperforms all the baselines at different channel conditions. This is a strong indication that maximizing $I(\hat{Z};Y)$ through contrastive learning yields better feature extraction and classification performance, even with undesirable and high noise conditions. As seen in Table V, CLAD outperforms both DeepJSCC and VIB, achieving the highest classification accuracy of 98.42% at SNR = 12 dB. This represents a 0.46% improvement over DeepJSCC (97.96%) and a 0.52% improvement over VIB at $\beta=0.001$ (97.90%). The higher accuracy of CLAD can be attributed to its ability to learn a more disentangled and task-relevant latent space using contrastive learning and adversarial disentanglement.
From Table V, CLAD achieves a very low IRI value of 0.039, drastically outperforming DeepJSCC (0.608) and VIB at $\beta=0.001$ (0.1931). This substantial difference shows that CLAD is effective at disentangling and removing task-irrelevant information from its encoded representations. This is further supported by the significantly lower adversarial accuracy of CLAD compared to DeepJSCC and VIB at $\beta=0.001$. From Fig. LABEL:fig:colored_mnist_reconstructions, we can visually see that image reconstructions from CLAD excel in retaining task-relevant information and preserving the core structure of the digits while discarding irrelevant information such as background color and other style features such as thickness or slant in the digit. This contrasts with DeepJSCCC and VIB, which retain unnecessary information such as background color and thickness and slant of the digit.
Interestingly, from Table V, VIB at lower $\beta$ values (e.g., $\beta=0.01$ and $\beta=0.1$) achieves IRI values that are closer to CLAD. However, the major drawback of VIB is that, while it improves privacy, informativeness and minimality, it comes at the cost of accuracy, especially as $\beta$ increases, as shown in the table. For example, at $\beta=0.1$, VIB achieves even lower IRI, however its classification accuracy drops to 93.29%, which is significantly lower than both CLAD and DeepJSCC. This demonstrates that CLAD maintains task performance while better preserving accuracy compared to VIB.
Figure 6: Reconstructed images from the received feature representations under different methods.